### v1

@@ Abstract
Reinforcement Learning (RL) consists of designing agents that make intelligent decisions without human supervision. When used alongside function approximators such as Neural Networks (NNs), RL is capable of solving extremely complex problems. Deep Q-Learning, a RL algorithm that uses Deep NNs, achieved super-human performance in some specific tasks. Nonetheless, it is also possible to use Variational Quantum Circuits (VQCs) as function approximators in RL algorithms. This work empirically studies the performance and trainability of such VQC-based Deep Q-Learning models in classic control benchmark environments. More specifically, we research how data re-uploading affects both these metrics. We show that the magnitude and the variance of the gradients of these models remain substantial throughout training due to the moving targets of Deep Q-Learning. Moreover, we empirically show that increasing the number of qubits does not lead to an exponential vanishing behavior of the magnitude and variance of the gradients for a PQC approximating a 2-design, unlike what was expected due to the Barren Plateau Phenomenon. This hints at the possibility of VQCs being specially adequate for being used as function approximators in such a context.
@@ 1.1 Context
VQCs are quantum circuits that depend on free parameters, which are iteratively updated by a classical optimizer to minimize an objective function estimated from measurements [4]. VQCs are especially suitable for NISQ devices, given their lower demands in the number of qubits and circuit depth, which mitigates the effect of noise [3]. Consequently, integrating VQCs into machine learning algorithms has become an intriguing field of study. VQCs used in such a context have been extensively researched in both supervised [5, 6, 7, 8] and unsupervised [9, 10] QML. Quantum Reinforcement Learning, on the other hand, remains largely unexplored.
@@
Just like VQCs have found applications as function approximators in both supervised and unsupervised ML domains, they can be effectively utilized in the same capacity in RL scenarios, serving as alternatives to DNNs. These quantum variational approaches in the context of RL have started gaining traction only recently, as evidenced by several recent studies [20, 21, 22, 23, 24], which are further discussed on the next subsection.
@@ 1.2 Related Work
This paper aims to directly address these questions. Specifically, we seek to investigate the impact of data re-uploading on both the performance and trainability of VQC-based Deep Q-Learning models. Our approach involves empirically analyzing the average return of two different hardware-efficient VQCs in benchmark environments and the behavior of their gradient’s norm and variance. Additionaly, we also verify how these metrics are affected as the number of qubits increases.
@@ 1.3 Contributions
  • Empirically verified that the gradient’s variance of the quantum models does not vanish exponentially in the aforementioned environments. Furthermore, we verified that the gradient’s variance increases when data re-uploading is used.
@@ 3.3 Cost Function and Gradient-Based Learning
$Q(s,a)=\bra{0^{\otimes n}}U(s,{\boldsymbol{\theta}})^{\dagger}(s)O_{a}U(s,{\boldsymbol{\theta}})(s)\ket{0^{\otimes n}}$ (9)
@@ 3.4 Barren Plateau Phenomenon
Put in words, the variance of the gradient of hardware-efficient VQC-based cost functions decays exponentially with the number of qubits [27]. Furthermore, [28] demonstrated that the more expressive the ansatz, the lower the variance in the cost gradient. Consequently, the cost landscape is flatter, making these circuits extremely hard to train. Nonetheless, it is important to emphasize that [37] observed that the Barren Plateau Phenomenon highly depends on the locality of the observable employed. To solve this issue, several strategies have been proposed [38]. Nonetheless, the problem severely limits the applicability of hybrid classical/quantum algorithms to solve practical problems, since it erases a potential quantum advantage with a VQC [3].
@@ 4.3 Data Re-Uploading’s Effect on Performance
Figures 2(a) and 2(b) show the considerable effects of data re-uploading and trainable input and output scaling on the performance of VQC-based Deep Q-Learning agents in the CartPole-v0 and Acrobot-v1 environments. Agents that do not use trainable output scaling perform very poorly, no better than random guessing, demonstrating the importance of matching the range of the Q-values estimated by the quantum circuit to the range of the optimal Q-values of the specific environment. While it is possible to achieve a considerable level of performance by multiplying the expectation values by a fixed value instead of using a classical trainable weight, [22] show that such a technique depends on the speed of convergence of the models and does not achieve the same level of performance as models that use trainable output scaling.
@@ 4.4 Data Re-Uploading’s Effect on Trainability
In this section, we turn our attentions to the trainability of the models from Figures 2(a) and 2(b). Since the models without trainable output scaling achieved a very poor level of performance, they were excluded from this analysis. The results can be seen in Figures 3(a) and 3(b).
Figure 4: Trainability of the models from Figures 2(a) and 2(b) with trainable output scaling in the CartPole-v0 (see Subfigure 3(a)) and Acrobot-v1 (see Subfigure 2(b)) environments. In both Subfigures, the left graph represents the gradient’s norm throughout training and the right graph the variance of the norm.
The data re-uploading models are noteworthy in several ways. Despite their greater circuit depth and enhanced expressivity, they exhibit the highest variance in the gradient norm among all models. This promising finding is in stark contrast to our initial expectations derived from [28], which suggests a trade-off between expressivity and trainability. In particular, Barren Plateaus are to be expected when using hardware-efficient ansatzes with high expressivities. However, in this case, the most expressive models are also the one with the highest gradients’ magnitude and variance. This is particularly surprising in the Cartpole-v0 environment, since the difference in the gradients’ norm and its variance between the data re-uploading model with input scaling and all the other models is substantial. In the next section, we study one of the possible causes for this behavior.
@@ 4.5 Trade-off Between Moving Targets and Gradient Magnitude
Figures 4(a) and 4(b) show how data re-uploading models perform for different values of $C$ and the behavior of their loss functions in the CartPole-v0 and Acrobot-v1 environments.
@@
This raises an interesting question. It is known that hardware-efficient VQCs suffer from the Barren Plateau Phenomenon [27], such that the variance of the gradients decays exponentially with system size. However, it appears that Deep Q-Learning is an inherently unstable algorithm. Intriguingly, certain hyperparameter configurations seem to allow Deep Q-Learning models to effectively learn good policies while sustaining considerable gradient magnitudes and variances. This hints at a potential advantage: might this inherent instability help counteract the Barren Plateau Phenomenon? In other words, could the utilization of highly expressive VQCs (such as data re-uploading ones) in such a context be somewhat resistant to Barren Plateaus? To verify these claims, it becomes imperative to research the gradient’s behavior for an increasing number of qubits. The most appropriate architecture for such a study is the UQC, since it allows the encoding of an input vector into an arbitrary number of qubits, leading to a higher flexibility in the choice of system size. Thus, in the next section we introduce this architecture and, then, in section 4.7, we analyze how the gradients behave as the number of qubits increases.
@@ 4.6 The Single and Multi-Qubit UQC
The reason for using these two types of encoding is two-fold. On the one hand, these two techniques allow us to study the impact of introducing entanglement on the performance of the models in the CartPole-v0 environment. On the other hand, the Full Encoding technique allows us to encode a given input vector in an arbitrary number of qubits that may even be greater than the number of features of the input vector. Thus, we may study how the performance and trainability behave as the number of qubits increases, see Section 4.7. Figures 6(a) and 6(b) show how each of these data encoding methods perform with and without entanglement for $2$ and $4$ qubits in the Cartpole-v0 and Acrobot-v1 environments, see Subfigures 6(a) and 6(b), respectively.
@@ 4.7 Gradient Behavior for Increasing System Sizes
One of the advantages of the multi-qubit UQC architecture is that one may encode any input vector into an arbitrary number of qubits. In particular, it is possible to increase the number of qubits even further than the number of features of the input vector. To test how the norm of the gradients and its variance change as the number of qubits increases, we trained the full encoding multi-qubit UQC model with the set of even numbers of qubits from $2$ to $12$ in the CartPole-v0 and Acrobot-v1 environments, see Figures 7(a) and 7(b), respectively.
@@
Both the magnitude of the gradient and its variance behave similarly during training for all models, while also achieving a similar range of values. However, it is important to note that there seems to be a decrease in the variance of the gradients for the Acrobot-v1 environment as the number of qubits increases, but this seems to be most noticeable after $10000$ training steps, after which most of the training should be over. These results are particularly interesting because they confirm the suspicions that arose throughout this work. The Barren Plateau Phenomenon states that the gradients decay exponentially with system size. However, for the VQC-based Deep Q-Learning algorithm, using the multi-qubit UQC model in the CartPole-v0 and Acrobot-v1 environments with the MSE loss function, that doesn’t seem to be the case. This hints at the possibility of hardware-efficient VQCs being especially suitable to be used as function approximators in Deep Q-learning since they might be resistant to the Barren Plateau Phenomenon, to some extent, due to the moving targets’ effects on the gradients of this algorithm.
However, only circuits that approximate $2$-designs are susceptible to Barren Plateaus [28]. To verify whether a circuit approximates a $2$-design, one can use the Hilbert-Schmidt norm of $A$ defined as:
[…]
where the latter integral is taken over all states according to uniformly sampling the parameters $\theta$ of our VQC, a methodology in line with the one used in [40], although oversimplified. The lower the value of the norm of $A$, the better it approximates a $2$-design. Table 1 shows values of the norm of $A$ for different circuits. Although oversimplified, it shows that the multi-qubit UQC indeed appears to approximate a $2$-design circuit.
Table 1: Values of the Hilbert-Schmidt norm of $A$ for different circuits
Circuit Value
Haar integral (sanity check) 0.022
Multi-qubit UQC with $4$ qubits and $5$ layers 0.067
$4$-qubits empty VQC 0.99
[…]
@@ 5 Conclusion
However, a concern would be that the increase in expressivity and circuit depth that arises from the use of data re-uploading could decrease the trainability of these models due to the Barren Plateau Phenomenon [27]. To investigate such a concern, we analyzed the norm of the gradients and the variance of this norm. From this analysis, we verified that the gradients achieve substantial values and actually increase when data re-uploading is used versus when it is not used. Furthermore, we empirically showed that this increase is due to an instability that is inherent to Deep Q-Learning, which is the fact that the targets are non-stationary. Alongside the Mean Squared Error loss function, this leads to a very unstable algorithm with substantial gradients. Moreover, we also verified that increasing the number of qubits of the multi-qubit UQC in the CartPole-v0 and Acrobot-v1 environments does not lead to a decrease in the gradients’ magnitude or variance.
6 Future Work

### v2

@@ Abstract
Reinforcement Learning (RL) consists of designing agents that make intelligent decisions without human supervision. When used alongside function approximators such as Neural Networks (NNs), RL is capable of solving extremely complex problems. Deep Q-Learning, a RL algorithm that uses Deep NNs, has been shown to achieve super-human performance in game-related tasks. Nonetheless, it is also possible to use Variational Quantum Circuits (VQCs) as function approximators in RL algorithms. This work empirically studies the performance and trainability of such VQC-based Deep Q-Learning models in classic control benchmark environments. More specifically, we research how data re-uploading affects both these metrics. We show that the magnitude and the variance of the model’s gradients remain substantial throughout training even as the number of qubits increases. In fact, both increase considerably in the training’s early stages, when the agent needs to learn the most. They decrease later in the training, when the agent should have done most of the learning and started converging to a policy. Thus, even if the probability of being initialized in a Barren Plateau increases exponentially with system size for Hardware-Efficient ansatzes, these results indicate that the VQC-based Deep Q-Learning models may still be able to find large gradients throughout training, allowing for learning.
@@ 1.1 Context
VQCs are quantum circuits that depend on free parameters, which are iteratively updated by a classical optimizer to minimize an objective function estimated from measurements [4]. VQCs are especially suitable for NISQ devices, given their lower demands in the number of qubits and circuit depth, which mitigates the effect of noise [3]. Consequently, integrating VQCs into machine learning algorithms has become an intriguing field of study. VQCs used in such a context have been extensively researched in both supervised [5, 6, 7, 8] and unsupervised [9, 10] QML.
@@
Just like VQCs have found applications as function approximators in both supervised and unsupervised ML domains, they can be effectively utilized in the same capacity in RL scenarios, serving as alternatives to DNNs. These quantum variational approaches in the context of RL have started gaining traction only recently, as evidenced by several recent studies [20, 21, 22, 23, 24, 25], which are further discussed on the next subsection.
@@ 1.2 Related Work
This paper aims to directly address these questions. Specifically, we seek to investigate the impact of data re-uploading on both the performance and trainability of VQC-based Deep Q-Learning models. Our approach involves empirically analyzing the average return of two different hardware-efficient VQCs in benchmark environments and the behavior of their gradient’s norm and variance. Additionaly, we also verify how these metrics are affected by an increasing system size.
@@ 1.3 Contributions
  • Empirically verified that the gradient’s variance of the quantum models increases in the early stages of training in the aforementioned environments. Furthermore, we verified that the gradient’s variance increases when data re-uploading is used.
@@ 3.3 Cost Function and Gradient-Based Learning
$Q(s,a)=\bra{0^{\otimes n}}U(s,{\boldsymbol{\theta}})^{\dagger}O_{a}U(s,{\boldsymbol{\theta}})\ket{0^{\otimes n}}$ (9)
@@ 3.4 VQC’s Trainability
Put in words, the variance of the gradient of hardware-efficient VQC-based cost functions decays exponentially with the number of qubits [28]. Furthermore, [29] demonstrated that the more expressive the ansatz, the lower the variance in the cost gradient. Consequently, the cost landscape is flatter, making these circuits extremely hard to train. Nonetheless, it is important to emphasize that [38] observed that the Barren Plateau Phenomenon highly depends on the locality of the observable employed. To mitigate this issue, several methods have been proposed, such as Barren-Plateau Free ansatzes [39, 40]. However, Cerezo et al. [41] claim that circuits that can be proved to be free from Barren Plateaus actually end up encoding the problem in a polynomial sub-space of the Hilbert Space, one that is classically simulatable.
It is important to note that the BP Phenomenon is a statement over the landscape on average [42], thus it does not exclude the possibility of there being fertile valleys - regions with high gradients near good local minima. Reasearch on warm starts aims to initialize the VQCs in such regions of the landscape [43, 44, 45]. Care must be taken, however, since regions with large gradients are not necessarily near a good local minima. In fact, it has been shown that hardware efficient VQCs’ landscapes are swamped with bad local minima [46].
@@ 4 Numerical Results
In this section, we present and analyse experimental numerical results concerning both the performance and trainability of VQC-based DQN agents. The results are summarized in table 1.
Table 1:
Figure Explanation Conclusions
Fig. 3 Analysis of the effect data re-uploading and input/output scaling have on the performance of QRL agents in CartPole-v0 and Acrobot-v1 environments. Data re-uploading, input scaling, and output scaling all increase the agents’ performance.
Fig. 4 Analysis of the trainability of the agents from Fig 3. Gradients tend to increase throughout training and Data Re-uploading models exhibit higher gradients than baseline models.
Figs. 5 and 6 Analysis of the loss and its gradients throughout training for different values of the target network update frequency. It’s the instability that arises from low target network update frequency values that leads to the gradients’ behavior observed in Fig. 4.
Fig. 7 Performance of the Single and Multi-Qubit UQC with and without entanglement on the Cartpole-v0 and Acrobot-v1 environments. When no entanglement is used, all the features of the state must be encoded into all qubits, otherwise the agent’s lack the necessary information to achieve (near) optimal policies.
Fig. 8 Variance of the gradient for $1000$ uniformly initialized UQC agents with local and global cost functions as the number of qubits increases. The variance of the gradient decays exponentially when a global cost function is used and in the border of a polynomial to exponential regime when a local cost function is used.
Fig. 9 Performance and norm/variance of the gradients throughout training of the UQC models for an increasing number of qubits. The gradients exhibit a similar behavior to what was previously seen. Moreover, as the number of qubits increases, these gradients remain high.
Figs. 10 and 11 Performance and Gradients of the UQC and Skolik models trained on a binary classification dataset for an increasing number of qubits. The UQC model consistently achieves higher accuracies than the Skolik models. Moreover, these accuracies do not decrease as the number of qubits increases. Finally, the gradients do not exhibit the same behavior as seen in Fig 9. Thus, the moving targets of Deep Q-Learning contribute to the large gradients found throughout training.
For the Skolik models, the input and output scaling weights were initialized as $1$s while the rotational parameters $\theta\sim U[0,\pi]$. For the UQC models, the weights $\vec{w}\sim\mathcal{N}(0,0.01)$ , the bias $\vec{b}$ were initialized as 00s, the rotational parameters $\varphi\sim U[0,\pi]$ and the output scaling weights as $1$s. For the full hyperparameters for each experiment, see A.
@@ 4.3 Data Re-Uploading’s Effect on Performance
Figures 3(a) and 3(b) show the considerable effects of data re-uploading and trainable input and output scaling on the performance of VQC-based Deep Q-Learning agents in the CartPole-v0 and Acrobot-v1 environments. Agents that do not use trainable output scaling perform very poorly, no better than random guessing, demonstrating the importance of matching the range of the Q-values estimated by the quantum circuit to the range of the optimal Q-values of the specific environment. While it is possible to achieve a considerable level of performance by multiplying the expectation values by a fixed value instead of using a classical trainable weight, [22] show that such a technique depends on the speed of convergence of the models and does not achieve the same level of performance as models that use trainable output scaling.
@@ 4.4 Data Re-Uploading’s Effect on Trainability
In this section, we turn our attentions to the trainability of the models from Figures 3(a) and 3(b). Since the models without trainable output scaling achieved a poor level of performance, they were excluded from this analysis. The results can be seen in Figures 4(a) and 4(b).
Figure 4: Trainability of the models from Figures 3(a) and 3(b) with trainable output scaling in the CartPole-v0 (see Subfigure 4(a)) and Acrobot-v1 (see Subfigure 4(b)) environments. In both Subfigures, the left graph represents the gradient’s norm throughout training and the right graph the variance of the norm.
These results are noteworthy in several ways. First, there is a pattern in the gradient’s norm and variance. From Subfigure 4(a), one can see that they increase in the training’s early stages, peak and then decrease on Cartpole-v0. Interestingly, they increase when the agent needs to learn the most (when it is initialized and the environment is new) and start to decrease around the $10000$ training-step mark when the agent should have done most of the learning already and started converging to a policy. However, the gradients of the models trained on Acrobot-v1 consistently increase throughout training, see Subfigure 4(b). This could possibly be explained by the fact that the agents never reached a consistently high return in this environment.
Moreover, despite their greater circuit depth and enhanced expressivity, the data re-uploading models exhibit the highest gradient’s variance among all models. This findind is in stark contrast to our initial expectations derived from [29], which suggests a trade-off between expressivity and trainability. We should nonetheless note that [29] claims that more expressive models are harder to train in the context of Barren Plateaus, that is, the gradient’s variance of such models decreases exponentially with system size upon initialization (when parameters are initialized uniformly spanning the whole Hilbert Space). Thus, our results and theirs are not mutually exclusive in the sense that our models may still suffer from the same problem upon uniform initialization covering the whole Hilbert Space, but our initialization in the context of Deep Q-learning leads to large gradients throughout training, a similar concept to warm-starts.
@@ 4.5 Trade-off Between Moving Targets and Gradient Magnitude
Figures 5(a) and 5(b) show how data re-uploading models perform for different values of $C$ and the behavior of their loss functions in the CartPole-v0 and Acrobot-v1 environments.
@@
This raises an interesting question. It is well known that hardware-efficient VQCs suffer from various trainability issues, from optimization landscapes swamped with bad local minima [47] to the Barren Plateau Phenomenon [28], such that the variance of the gradients decays exponentially with system size. However, it appears that Deep Q-Learning is an inherently unstable algorithm due to the moving targets. Intriguingly, certain hyperparameter configurations seem to allow Deep Q-Learning models to effectively learn good policies while sustaining considerable gradient magnitudes and variances. This hints at a potential advantage: might this inherent instability help counteract the aforementioned trainability issues? We have seen that the gradient’s variance of the quantum models increases throughout training. Thus, even if these same models suffer from the Barren Plateau Phenomenon, it may be possible to train them in the context of Deep Q-Learning due to the large gradients seen throughout training. To verify these claims, it becomes imperative to research the gradient’s behavior for an increasing number of qubits. The most appropriate architecture for such a study is the UQC, since it allows the encoding of an input vector into an arbitrary number of qubits, leading to a higher flexibility in the choice of system size. Thus, in the next section we introduce this architecture and, then, in section 4.7, we analyze how the gradients behave as the number of qubits increases.
@@ 4.6 The Single and Multi-Qubit UQC
The reason for using these two types of encoding is two-fold. On the one hand, these two techniques allow us to study the impact of introducing entanglement on the performance of the models in the CartPole-v0 environment. On the other hand, the Full Encoding technique allows us to encode a given input vector in an arbitrary number of qubits that may even be greater than the number of features of the input vector. Thus, we may study how the performance and trainability behave as the number of qubits increases, see Section 4.7. Figures 7(a) and 7(b) show how each of these data encoding methods perform with and without entanglement for $2$ and $4$ qubits in the Cartpole-v0 and Acrobot-v1 environments, see Subfigures 7(a) and 7(b), respectively.
@@ 4.7 Gradient Behavior for Increasing System Sizes
One of the advantages of the multi-qubit UQC architecture is that one may encode any input vector into an arbitrary number of qubits. In particular, it is possible to increase the number of qubits even further than the number of features of the input vector.
We first start by analysing the gradient’s variance of the multi-qubit UQC as the number of qubits increases when a uniform initialization that spans the whole Hilbert Space is used. Thus, we only consider the first training step of CartPole-v0 and plot the variance of the norm of the gradient of $1000$ agents for every even number of qubits between $2$ and $16$, see Figure 8.
Figure 8: Variance of the Gradient’s Norm across $1000$ uniformly sampled parameters spanning the entire Hilbert Space. Two cost functions are tested, a global cost function and a local cost function. For the full set of hyperparameters, see 7.
When a global cost function is used, the variance decays exponentially with the number of qubits. However, for a local cost function (the one used throughout this work for the QRL agents), it seems to decay in a regime on the border of polynomial to exponential (as indicated by the R2 error on the fit), a result consistent with [38].
However, we are particularly interested to see how the gradients behave throughout training when using fine-tuned initialization strategies. Thus, we trained the full encoding multi-qubit UQC model with the set of even numbers of qubits from $2$ to $12$ in the CartPole-v0 and Acrobot-v1 environments, see Figures 9(a) and 9(b), respectively.
Both the magnitude of the gradient and its variance behave similarly during training for all models, while also achieving a similar range of values. However, it is important to note that there seems to be a decrease in the variance of the gradients for the Acrobot-v1 environment as the number of qubits increases, but this seems to be most noticeable after $10000$ training steps, after which most of the training should be over. These results are particularly interesting because they confirm the suspicions that arose throughout this work. Throughout training, all the models exhibit large gradients even as the number of qubits increases. This seems to be similar to warm-starts, since even though these landscapes are characterized by exponentially vanishing gradients, the models are able to navigate the landscape in zones with high gradients, facilitating training. This hints at the possibility of hardware-efficient VQCs being especially suitable to be used as function approximators in Deep Q-learning.
@@
4.8 Gradient Behavior in Supervised Learning
So far, we have claimed that the high gradients found throughout training by the VQC-based Deep Q-Learning models are due to the instability introduced by the moving targets. To verify whether this is true, we trained both the Skolik Data Re-uploading and the Full Encoding Multi-Qubit UQC on the more common task of binary classification for an increasing number of qubits. Since the Skolik Architecture requires a number of qubits equal to the number of features of the datapoints, we used scipy’s make_classification method to generate a dataset of $500$ datapoints for each number of qubits tested. To keep this as close to the original RL scenario as possible, we also used the MSE cost function and the same observables as those used previously. The accuracies achieved on the training and validation datasets as well as the MSE cost function throughout training can be seen in Fig 10
Interestingly, the Multi-Qubit UQC model consistently outperforms the Skolik model, achieving higher training and validation accuracies and a lower cost value. Moreover, this performance does not decrease as the number of qubits increases, unlike what happens to the Skolik Model, where for $12$ qubits the validation accuracy does not even reach $70\%$. However, more than the performance of such models in terms of accuracy, we want to know how the gradients behave throughout training for a comparison with what we have seen in the VQC-based Deep Q-Learning agents, see Figure 11.
The Skolik Models showcase a clear trend in both the norm and the variance of the gradients. As the number of qubits increases, both metrics decrease upon initialization and behave similarly throughout training: they first decrease and then converge. However, for the UQC models, there is not a clear trend in the behavior of the norm of the gradient. Moreover, while the variances across the models with different numbers of qubits behave similarly, there is not a clear decrease in the variance unlike what is seen for the Skolik models.
These results show that the gradients behave differently in the context of supervised learning when compared to Deep Q-learning. This observation confers additional support to the hypothesis that it is the instability caused by the moving targets that allow for high gradients throughout training.
Figure 10: Training and Validation accuracy as well as MSE cost function throughout training for the Skolik Data Re-Uploading (Subfigure 10(a)) and Full-Encoding Multi-Qubit UQC (see Subfigure 10(b)) trained on a binary classification problem. The full set of hyperparameters can be seen in 8.
Figure 11: Norm and Variance of the norm of the gradient vector throughout training for the Skolik Data Re-Uploading (Subfigure 10(a)) and Full-Encoding Multi-Qubit UQC (see Subfigure 10(b)) trained on a binary classification problem. The full set of hyperparameters can be seen in 8.
@@ 5 Conclusion
However, a concern would be that the increase in expressivity and circuit depth that arises from the use of data re-uploading could decrease the trainability of these models. To investigate such a concern, we analyzed the norm of the gradients and the variance of this norm. From this analysis, we verified that the gradients achieve substantial values and actually increase when data re-uploading is used versus when it is not used. Furthermore, we empirically showed that this increase is due to an instability that is inherent to Deep Q-Learning, which is the fact that the targets are non-stationary. Alongside the Mean Squared Error loss function, this leads to a very unstable algorithm with substantial gradients. Moreover, we also verified that increasing the number of qubits of the multi-qubit UQC in the CartPole-v0 and Acrobot-v1 environments does not lead to a decrease in the gradients’ magnitude or variance thorughout training.
5.1 Future Work